# Uncertainty-driven Adaptive Exploration

### v2

@@ 2.2 Generality of the ADEU framework
We now elaborate on the fact that ADEU is a generic adaptive exploration framework. First, it can readily use as $f(s)$ any predefined mechanism—e.g., some epistemic uncertainty-measuring mechanism, or mechanisms used in intrinsic motivation methods to approximate visitation frequency. Moreover, it is a framework that can emulate most, if not all, existing adaptive exploration techniques. This can be achieved by designing an as-complex-as-required $f(s,\cdot)$— e.g., an $f(s,\cdot)$ mechanism that outputs a value considering previous states or actions, or other constraints.
@@
To see this, consider first, the intrinsic motivation (IM) family, which rewards the agent whenever a novel state is visited. This non-stationary reward, termed as $r_{\text{intrinsic}}(s)$, is added to the external reward provided by the environment, as described in Equation 2. Then, the agent acts greedily to maximize 2.
@@
To calculate the $r_{\text{intrinsic}}(s)$, IM-oriented methods [1, 3, 17, 18] propose numerous mechanisms that can quantify the novelty of a state. For example, RND [3] proposes to calculate the visitation frequency as the approximation error between two networks named Target and Predictor. Predictor network is trained to predict the Target’s output for a given state. The difference between the output of the two networks (Equation 3) approximates the visitation frequency for that specific state.
@@
ADEU can utilize this approximation error as an uncertainty-measuring mechanism. That is, an instance of ADEU that defines uncertainty as the visitation frequency of a state can utilize $f(s)=r^{\text{RND}}_{\text{intrinsic}}(s)$. Of course, any other mechanism approximating visitation frequency, such as the ones proposed in other IM methods [1, 17, 26] can also be employed.
One can also create an instance of ADEU that utilizes an uncertainty mechanism similar to the ones defined in epistemic-uncertainty driven methods [4, 13]. Epistemic uncertainty-driven methods employ different representations of the same Q-function to measure how certain an agent is about its policy in a specific state. For example, the UCB [4, 13] strategy greedily selects actions that maximize both the $Q_{\text{mean}}$ and the $\lambda Q_{\text{std}}(s,a)$ as shown in Equation 4.
@@
where $Q_{\text{mean}}(s,a)$ is the mean $Q$-value calculated across the different representations of the $Q$-function, $Q_{\text{std}}(s,a)$ is the corresponding standard deviation, and $\lambda$ is a positive constant. Given this, we can also create an ADEU instance that mirrors the above uncertainty-measuring method, via setting $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$. Table 1 summarizes the discussion above.
Table 1: Instances of ADEU using different uncertainty-measuring mechanisms—examples only; any IM mechanism and any mechanism calculating confidence over $\pi$ can be used.
@@
Uncertainty is calculated regarding the novelty of each state $f_{\text{RND}}(s)=\left\lVert f_{\text{predictor}}(s)-f_{\text{target}}(s)\right\rVert_{2}^{2}$
Uncertainty is calculated as the confidence of an agent about its policy $f_{\text{UCB}}(s)=\lambda Q_{\text{std}}(s,\pi(s))$
@@
Table 2 presents certain key adaptive exploration works, and their instantiation within the ADEU framework. See also Section 4 for more details.
@@
Table 2: Previous adaptive exploration works and how these can be instantiated in ADEU.
Previous Work Description Special Case of ADEU
Value-Difference Based Exploration[29] Uses $\epsilon$-greedy. Modifies $\epsilon$ at every state wrt. TD-Error A Bernoulli distr. can be used to simulate $\epsilon$-greedy. f(s)f(s)=TD-Error(s)
$\epsilon z$-greedy [6] Agent performs an “option” (i.e, a sequence of actions) with constant probability $\epsilon$. The length of that option is sampled from a heavy-tailed distribution. With prob. $1-\epsilon$, agent uses its policy $\pi(s)$. We will use a Bernoulli to simulate whether an agent will act according to its $\pi(s)$ or not. Then $f(s,n)$, where $n$ is sampled from the heavy-tailed distribution at the first call of $f(s,)$, will change Bernoulli’s pp. [^6] At the first call, $f(s,n)$ where $n$ is an optional argument not used in the first call, it will sample the $n$. Then, at every next call of the $f(s,n)$ the $n$ parameter will be used as an argument and decreased by $1$.
Pislar et al. [23] This is a general work discussing ‘when should agents explore’. It proposes intra-episodic exploration after an agent receives a signal. This exploration ‘session’ lasts for a fixed number of timesteps. A Bernoulli is required to alternate (by modifying $p$) between $\pi(s)$ and a random action. $p$ will be modified according to $f(s)$, which will replace the signals. After $f(s)$ changes $p$ to promote exploration, it will output a constant value for a fixed number of timesteps.
Goal-based exploration (“Go-explore” approach) [8, 9, 30] An agent first replicates a trajectory (or a ‘return’ or ‘go’ phase), and then explores. Either a Bernoulli or a Gaussian can be used, depending on the action space. $f(s)$ will be a binary function.
@@
Table 2 explains how ADEU incorporates previous works on adaptive exploration. Given that ADEU can subsume existing adaptive exploration methods as special cases, it is reasonable to expect that it can replicate their performance in the tested domains. Thus, ADEU is well-suited for tasks with both dense and sparse reward structures. [^7] Regardless, we plan to test ADEU in environments with sparse reward functions also, like ATARI2600 [2]. In addition, by using more complex definitions of “uncertainty” (e.g., Nikolov et al. [20], Badia et al. [1]) we can enhance exploration in domains with less informative and sparse reward functions.
Motivation example cont’d:
@@
We now continue with the technical setup and the results of the motivation example. We design a large-scale grid environment of size $1500\times 1500$, extending the setup shown in Figure 1. In this environment, we evaluate four different agents: (i) a standard $Q$-learning agent using $\epsilon$-greedy exploration, (ii) an intrinsic motivation (IM) agent built on top of $Q$-learning, (iii) a UCB-based ensemble agent composed of multiple $Q$-learners, and (iv) an instance of ADEU. For the IM agent, we adopt ideas from [26] without the need for approximation due to the discrete state space. The ADEU agent adjusts its action selection based on state visitation frequency, favoring exploration in less-visited states and exploiting learned values in frequently visited ones. It uses the same mechanism as the IM agent. Additional implementation details about the environment, the probabilistic distribution and activation function choices are provided in the Appendix.
@@
Table 3 shows the maximum reward and the mean reward during learning, computed as follows. We ran $5$ different runs (seeds) for all agents, and the following process was performed for each of the runs: Once every $100$ episodes, an “evaluation” episode was executed, in which the agent was using greedily its policy (i.e., no actions were selected for exploration during that episode); and we stored the reward accumulated during that episode, and all such episodes in each run. We then create an “average run” by averaging over these episodes’ values across all runs; and compute the mean episodic reward and the maximum episodic reward in that average run. Note that the maximum episodic reward was achieved after convergence.
@@
Table 3: Mean and Max episodic rewards of different exploration strategies in the Frozen Lake domain. Results are averages over 55 runs. ADEU was built on top of a $Q$-learning agent.
@@
$\epsilon$-greedy 334.0; 983.0
ADEU (Using visitation frequency) 88,854.0; 104,487.0
IM 393.0; 989.0
UCB 36,070.0; 104,487.0
As shown in Table 3, ADEU demonstrates superior performance compared to the other methods by rapidly learning a policy that consistently guides it toward the target price. Specifically, the ADEU agent avoids excessive exploration of frequently visited states—those characterized by low uncertainty according to the employed function $f(s)$. Instead, it tends to replicate the learnt trajectory until completion before initiating further exploration; and manages to learn the correct trajectory. By contrast, both the $\epsilon$-greedy and IM algorithms fail to learn a successful trajectory to the goal, likely due to their persistent over-explorative behavior. Finally, UCB learns the correct trajectory, but it requires nearly four times as many episodes as ADEU (16,800 vs 4,515 for ADEU). Its average reward is approximately half that of ADEU’s. Note that in our main experiments involving more complex action and state spaces, UCB always ranks behind at least one ADEU variant.
@@
Figure 2 (top figure) shows both the actual policy $\pi(s)$ and the action selected from the ADEU agent (so, using Equation 1) during a single episode in our Frozen Lake environment. In addition, the bottom figure shows the value of $g(f(s))$ during the same episode. As expected, at the beginning of the episode $g(f(s))$ outputs low values since the agent visits states that it is certain about. Hence, the action selected by the ADEU agent coincides with the one dictated by $\pi(s)$. As the agent visits states with high uncertainty, it does not trust its so far learned policy and starts exploring them. That is, the action selected by ADEU and the policy $\pi(s)$ differ at the end of the episode, since $g(f(s))$ increases. We should also note that the action suggested by $\pi(s)$ in the later timesteps of the episode is not the ideal one. However, the action sampled for the exploration strategy was the correct one, allowing the agent to “survive” for more timesteps and gain useful experience. Thus, indeed, the exploration phase should have begun at the later timesteps; had it started earlier, the agent would not have “survived” to extend the known optimal policy for the early timesteps to the later ones.
ADEU’s variants for safe exploration
@@
Now, as discussed and seen in the example, ADEU avoids exploration in states with low uncertainty. Moreover, it can be extended so that it conducts safe exploration—i.e., to not explore in unsafe states [33, 34]. By simply defining a new mechanism that also considers this extra constraint, ADEU can avoid exploring these states.
@@
More precisely, suppose that a mentor has already informed the agent about some potential unsafe states (like in Thomas et al. [27]) and the $\pi(s)$ policy the agent should follow in these particular states. We would like the agent to not explore finding a better policy in these specific states. Hence, by designing an $f(s)$ that outputs a low uncertainty in these states, an ADEU agent can alternate between exploration and exploitation proportional to the safety of the current state.
@@
Similarly, the agent can be informed of the cost function of the environment (like in Tian et al. [28]). By augmenting $f(s)$ with the cost function, ADEU can combine adaptive exploration with safety. Even if the ADEU agent does not obtain any knowledge regarding the unsafe states or the cost function, it can interact with the environment to approximate it. Along with this “approximation” process, the agent can still respect the safety constraints.
3 Experimental Setup and Results

### v3

@@ 2.2 Generality of the ADEU framework
In this section, we elaborate on the fact that ADEU is a generic adaptive exploration framework. First, ADEU can readily use as $f(s)$ any predefined mechanism—e.g., some epistemic uncertainty-measuring mechanism, or mechanisms used in intrinsic motivation methods to approximate visitation frequency. Moreover, it is a framework that can emulate most, if not all, existing adaptive exploration techniques. This can be achieved by designing an as-complex-as-required $f(s,\cdot)$— e.g., an $f(s,\cdot)$ mechanism that outputs a value considering previous states or actions, or other constraints.
@@
To see this, consider first the intrinsic motivation (IM) family, which rewards the agent whenever a novel state is visited. This non-stationary reward, termed as $r_{\text{intrinsic}}(s)$, is added to the external reward provided by the environment, as described in Equation 2. Then, the agent acts greedily to maximize 2.
@@
To calculate the $r_{\text{intrinsic}}(s)$, IM-oriented methods [Badia2020NeverGiveUp, burda2018rnd, Lobel2023FlippingCoins, mahankali2024random] propose numerous mechanisms that can quantify the novelty of a state. For example, RND [burda2018rnd] proposes to calculate the visitation frequency as the approximation error between two networks named Target and Predictor. Predictor network is trained to predict the Target’s output for a given state. The difference between the output of the two networks (Equation 3) approximates the visitation frequency for that specific state.
@@
ADEU can utilize this approximation error as an uncertainty-measuring mechanism. That is, an instance of ADEU that defines uncertainty as the visitation frequency of a state can utilize $f(s)=r^{\text{RND}}_{\text{intrinsic}}(s)$. In a similar way, any other mechanism approximating visitation frequency, such as the ones proposed in other IM methods [Badia2020NeverGiveUp, Lobel2023FlippingCoins, tang2017exploration] can also be employed.
One can also create an instance of ADEU that utilizes an uncertainty mechanism similar to the ones defined in epistemic-uncertainty driven methods [Chen2017ucb, lee2021sunrise]. Epistemic uncertainty-driven methods employ different representations of the same Q-function to measure how certain an agent is about its policy in a specific state. For example, the UCB [Chen2017ucb, lee2021sunrise] strategy greedily selects actions that maximize the upper bound of the $Q$-function shown in Equation 4.
@@
where $Q_{\text{mean}}(s,a)$ is the mean $Q$-value calculated across the different representations of the $Q$-function, $Q_{\text{std}}(s,a)$ is the corresponding standard deviation, and $\lambda$ is a positive constant. Given this, we can also create an ADEU instance that utilizes the above uncertainty-measuring mechanism by setting $f(s)=\lambda Q_{\text{std}}(s,\pi(s))$. Table 1 summarizes the discussion above.
Table 1: Instances of ADEU using different uncertainty-measuring mechanisms—examples only; any IM or mechanism calculating confidence over $\pi$ can be used.
@@
Novelty of $s$ $f_{\text{RND}}(s)=\left\lVert f_{\text{predictor}}(s)-f_{\text{target}}(s)\right\rVert_{2}^{2}$
Confidence on $\pi(s)$ $f_{\text{UCB}}(s)=\lambda Q_{\text{std}}(s,\pi(s))$
@@
Table 2 presents certain key adaptive exploration works, and their instantiation within the ADEU framework. See also “Related Work” for more details.
@@
Table 2: Previous adaptive exploration methods as special cases of ADEU.
Description Special Case of ADEU
tokic2010adaptive: $\epsilon$-greedy where $\epsilon$ is adapted at each state based on TD error. Utilizing a Bernoulli distribution as $D$. Let $f(s)$ = TD-Error(s)(s).
dabney2021ez_greedy: Executes a fixed-length option with prob. $\epsilon$; otherwise follows $\pi(s)$. Length sampled from a heavy-tailed distribution. A Bernoulli decides between $\pi(s)$ and an option. We will utilize a more complex uncertainty-measuring mechanism $f(s,n)$, that calculates uncertainty according (also) to a parameter nn—this is still allowed by ADEU as previously discussed. This mechanism sets the probabilities $p$ of the Bernoulli based on $n$, which is sampled once and decremented at each step.666nn is sampled in the first call to $f(s,n)$, then decremented in subsequent calls.
pislar2022when: Explores intra-episodically after a trigger signal, for a fixed duration. $f(s)$ replaces the trigger. A Bernoulli distribution is utilized. $f(s)$ modifies the probability $p$ to favor exploration, then outputs a constant $p$ for a fixed number of steps.
ecoffet2019goexplore, Ecoffet2021first_return_then_explore, torne2023breadcrumbs: The agent first returns to a state, then explores. Use Bernoulli or Gaussian. $f(s)$ is binary and switches phases.
Table 2 explains how ADEU incorporates previous works on adaptive exploration. For instance, ADEU can emulate  dabney2021ez_greedy by utilizing a Bernoulli distribution, which will “decide” over two different scenarios: (i) policy $\pi(s)$ with probability 1p1-p; and (ii) an option with probability $p$. The $p$ parameter of the Bernoulli distribution will dynamically change according to $f(s,n)$. In more detail, the initial value for $n$ is 0. At the first call of $f(s,n)$, $n$ will be sampled from a heavy-tail distribution and then, $p=f(s,n)$, where $f(s,n)=1$ when $n>0$, and otherwise it follows the procedure described in dabney2021ez_greedy. In addition, ADEU can also improve $\epsilon z$-greedy by incorporating more complex definitions of uncertainty, maintaining the use of the options. Hence, “extensions” of $\epsilon z$-greedy [biedenkapp2021temporl, Lee2024UTE], that improve dabney2021ez_greedy by modifying the way that the length $n$ of a trajectory is decided, can also be emulated by ADEU. In addition, instead of having a Bernoulli distribution, we can utilize a Categorical one. The process of modifying $p$ will remain the same, however, in this scenario, we can simplify the “task” of finding an option.
@@
Given that ADEU can subsume existing adaptive exploration methods as special cases, it is reasonable to expect that it can replicate their performance in the tested domains. Thus, ADEU is well-suited for tasks with both dense and sparse reward structures. For instance, ADEU can solve the extremely demanding DeepSea domain [osband2019deepSea] by emulating the dabney2021ez_greedy adaptive exploration method.
Motivating example cont’d:
@@
We now continue with the technical setup and the results of the motivation example. We design a large-scale grid environment of size $1500\times 1500$, extending the setup shown in Figure 1. In this environment, we evaluate four different agents: (i) a standard $Q$-learning agent using $\epsilon$-greedy exploration, (ii) an intrinsic motivation (IM) agent built on top of $Q$-learning, (iii) a UCB-based ensemble agent composed of multiple $Q$-learners, and (iv) an instance of ADEU. For the IM agent, we adopt ideas from [tang2017exploration] without the need for approximation due to the discrete state space. The ADEU agent adjusts its action selection based on state visitation frequency, favoring exploration in less-visited states and exploiting learned values in frequently visited ones. It uses the same mechanism as the IM agent. (See Appendix for more details.)
@@
Table 3 shows the maximum reward and the mean reward during learning, computed as follows. We employed $5$ different runs (seeds) for all agents, with each run comprising $3\times 10^{4}$ episodes. An episode terminates when: (i) the agent reaches the goal; or (ii) the agent falls in a hole; or (iii) the agent acts/survives $t>3\times 1500$. The following process was performed for each run: Once every $100$ episodes, an “evaluation” episode was executed, in which the agent greedily used its policy (i.e., no actions were selected for exploration during that episode); and we stored the reward accumulated during that episode, and all such episodes in each run. We then created an “average run” by averaging over these episodes’ values across all runs; and computed the mean episodic reward and the maximum episodic reward in that average run. Note that the maximum episodic reward was achieved after convergence.
@@
Table 3: Mean and Max episodic rewards in the Frozen Lake domain (average over 55 runs). ADEU was built on top of a $Q$-learning agent.
@@
$\epsilon$-greedy 334;    983
ADEU 88,854;   104,487
IM [tang2017exploration] 393;    989
UCB [Chen2017ucb] 36,070;   104,487
Figure 2: Environment used in testing safety. In this particular figure, the dimensions of the grid are $10\times 10$. The red crosses represent the unsafe states.
As shown in Table 3, ADEU demonstrates superior performance compared to the other methods by rapidly learning a policy that consistently guides it toward the target price. Specifically, the ADEU agent avoids excessive exploration of frequently visited states—those characterized by low uncertainty according to the employed function $f(s)$. Instead, it tends to replicate the learnt trajectory until completion before initiating further exploration; and manages to learn the correct trajectory. By contrast, both the $\epsilon$-greedy and IM algorithms fail to learn a successful trajectory to the goal, likely due to their persistent over-explorative behaviour. Finally, UCB learns the correct trajectory, but it requires nearly four times as many episodes as ADEU (16,800 vs 4,515 for ADEU). Its average reward is approximately half that of ADEU’s. Note that in our main experiments involving more complex action and state spaces, UCB always ranks behind both ADEU variants (see Table 5).
@@
Figure 3 (top subfigure) shows both the actual policy $\pi(s)$ and the action selected from the ADEU agent (so, using Equation 1) during a single episode in our Frozen Lake environment. In addition, the bottom subfigure shows the value of $g(f(s))$ during the same episode. As expected, at the beginning of the episode, $g(f(s))$ outputs low values since the agent visits states that it is certain about. Hence, the action selected by the ADEU agent coincides with that dictated by $\pi(s)$. As the agent visits states with high uncertainty, it does not trust its so far learned policy and starts exploring them. As of that, the action selected by ADEU and the policy $\pi(s)$ differ at the end of the episode, since $g(f(s))$ increases. Note also that the action suggested by $\pi(s)$ in the later timesteps of the episode is not correct. However, the action sampled for exploration was the correct one, allowing the agent to “survive” for more timesteps and gain useful experience. Thus, indeed, the exploration phase should have begun at the later timesteps; had it started earlier, the agent would not have “survived” to extend the known optimal policy.  [^7] We expect the same behaviour in more difficult multi-dimensional domains such as the ones described in Section 3. However, we cannot create a figure similar to Figure 3 for those, since both $\pi(s),a\in\mathbb{R}^{n}$, where $n>2$.
Generalizing Definition 2.1:
Definition 2.1, can be further generalized to also include sparse (but still increasing) reward settings. To show that, we will first define two types of states:
Definition 2.3 (Non-Terminal-Hyper-State).
A Non-Terminal-Hyper-State is defined as a set consisting of: (i) N1N-1 states $s_{n}:r(s_{n},a)=0\forall a$; and (ii) one state $s_{d}:\exists a_{d}\text{ where }r(s_{d},a_{d}\neq 0)$. All $s_{n}\text{ and }s_{d}$ are non-terminal states.
Definition 2.4 (Terminal-Hyper-States).
A Terminal-Hyper-State is defined as the set of consecutive terminal states.
Using the aforementioned definitions of Hyper-States, Definition 2.1 can be generalized also to include “Sparse Increasing Reward functions Single Agent Games” as follows:
Definition 2.5 (Sparse Increasing Reward Functions Single Agent games).
We define as “Sparse Increasing Reward Function Single Agent Game” a game that replaces states with Hyper-States in Definition 2.1. In more detail, $s^{*}$ states are replaced by Nom-Terminal-Hyper-State with $r(s_{d},a_{d})>0$, $s^{!}$ by Non-Terminal-Hyper-State with $r(s_{d},a_{d})<0$ and all the other states are replaced by Terminal-Hyper-States.
Corollary 2.2.
Theorem 1 holds for finding the best hyper-action in Sparse Increasing Reward Functions Single Agent Games. A hyper-action can be defined as a sequence of actions, exactly like “options”. The uncertainty-measuring mechanism in ADEU variances that act in games satisfying Definition 2.5 should calculate uncertainty in Hyper-States (i.e., a batch of non-terminal states).
2.3 ADEU for safe exploration
@@
Now, as discussed and seen in the example, ADEU avoids exploration in states with low uncertainty. Moreover, it can be extended so that it conducts safe exploration—i.e., to not explore in unsafe states [wachi2023safe, ijcai2022p3o]. By simply defining a new mechanism that also considers this extra constraint, ADEU can avoid exploring these states.
@@
More precisely, suppose that a mentor has already informed the agent about some potential unsafe states (like in thomas2021imagining_near_future) and the policy $\pi(s)$ the agent should follow in these particular states. We would like the agent to not explore finding a better policy in these specific states. Hence, by designing an $f(s)$ that outputs low uncertainty in these states, an ADEU agent can alternate between exploration and exploitation proportionally to the safety of the current state. Such an ADEU agent, will simply follow Equation 1.
@@
Alternatively, the agent can be informed about the cost function of the environment (like in tian2024reinforcement). By replacing $f(s)$ with the cost function, ADEU can combine adaptive exploration with safety. Even if the ADEU agent is not informed regarding the unsafe states or the cost function, it can interact with the environment to approximate it. Along with this “approximation” process, the agent can still respect the safety constraints. In that case, ADEU only requires an initial safe policy as in as2025actsafe. That is, Eq. 5 allows for the aforementioned “safety-driven” or “safety-aware” behaviour, without requiring knowledge of the cost function.
$a\sim\mathcal{D}\left(\pi(s),g(\nicefrac{{1}}{{f_{\text{safety}}(s)}})\right)$ (5)
where $f_{\text{safety}}(s)$ is the expectation of the environment’s cost function $\mathbb{E}_{s,a}\max_{a}\{g_{\text{cost}}(s,a)\}$ (where $g_{\text{cost}}(s,a)$ follows the standard definition described in [altman1999cmdp]), which is learned by interactions.
By sampling from Equation 5, the agent will replicate the known policy $\pi(s)$ in a state with high expected cost; in any other state, the agent will explore randomly. In more detail, in a state with high $f_{\text{safety}}(s)$, the sampled action $a$ will not deviate from $\pi(s)$, given the low variance in Equation 5. By contrast, in a state with low $f_{\text{safety}}(s)$, the sampled action will promote exploration. Finally, Equation 6 shows how we can combine Equations 1 and 5 to achieve both (i) safe exploration when required; and (ii) uncertainty-driven adaptive exploration in states where safety does not matter:
$a\sim\mathcal{D}\left(\pi(s),g(\nicefrac{{f_{\text{uncertainty}}(s)}}{{\lambda f_{\text{safety}}(s)}})\right)$ (6)
In Eq. 6 $f_{\text{uncertainty}}(s)$ is the aforementioned uncertainty-measuring mechanism and all other notation (except $\lambda$) follows Eq. 5. Note that $\lambda$ generally acts as a term (or mechanism) that ensures the safety term “dominates” when safety is critical. For example, $\lambda$ can take “conditioned” values for a state $s$ e.g., $\lambda=[max_{a}\{Q(s,a)\}$ if $\mathbb{E}_{s,a}\max_{a}\{g_{\text{cost}}(s,a)\}>max_{a}\{Q(s,a)\}$, otherwise 1 ]]. Therefore, the $f_{\text{safety}}(s)$ term will be amplified when the expected cost is high.
In this case, in states with high expected cost, the agent will act according to the known policy, in other states, it will act according to uncertainty-driven adaptive exploration. Thus, Equation 6 proposes a general agent that can take the best of both worlds.
Table 4: Performance metrics of different exploration strategies in the Safety Frozen Lake domain. Results are averages over 55 runs. ADEU was built on top of a $Q$-learning agent.
Exploration Strategy Mean & Max Reward Mean Constr.
$\epsilon$-greedy 210; 739 0.34
IM [tang2017exploration] 267 ; 780 0.55
UCB [Chen2017ucb] 12,605 ; 65,687 0.10
ADEU (using Equation 1; f(s)=f(s)=Visitation freq.) 59,865; 104,487 0.01
ADEU (using Equation 5, i.e., safety-aware) 7;7 0.00
ADEU (using Equation 6; $f_{\text{uncertainty}}=$Visitation freq.) 19,255; 85,087 0.00
Figure 2 shows a variant of the modified Frozen Lake that also incorporates some constraints. In more detail, the red crosses represent non-terminal states that the agent should not visit. In cases that the agent visits such states, a safety constraint is violated, and also the agent gets an immediate reward equal to 0. Otherwise, it is rewarded according to the previous example. In addition, we conducted experiments in a $1500\times 1500$ grid similarly to Section 2.2.
To test both aforementioned scenarios, we create $3$ different instances of ADEU: (i) an agent that has been pretrained on a safe policy in states neighboring unsafe states, and explores using the visitation frequency definition of uncertainty (Equation 1); (ii) a safe ADEU agent that has been pretrained on the same safe policy, but explores according only to the cost function (Equation 5), which is learned simultaneously with policy $\pi(s)$; and (iii) a safe ADEU agent built on top of the aforementioned one, which explores the environment according both to the learned cost function and the visitation frequency definition of uncertainty (Equation 6).
Table 4 shows the mean and maximum episodic evaluation rewards throughout the training process. In addition, it shows the mean episodic constraint violations. All ADEU instances outperform their opponents since they achieve fewer constraint violations. Since all algorithms tested in a domain requiring efficient exploration, the ADEU instance that only uses the learned cost function (Equation 5) did not yield a high reward due to the lack of (efficient) exploration in safe states. By contrast, the ADEU instance of Equation 1 effectively explores the environment and finds the final goal faster, but at the cost of some constraint violations—still violates an order of magnitude fewer constraints than its non-ADEU opponents. Finally, the ADEU instance that combines safety with uncertainty-driven adaptive exploration (Equation 6) yields a high reward, achieving 0 constraint violations, utilizing the best of both worlds (uncertainty-driven adaptive exploration and safe exploration).
These results provide interesting intuitions for ADEU potential for safe exploration. As discussed in Section 5, we are actively working towards theoretical guarantees for ADEU in safety domains.
3 Experimental Evaluation